### v1

# Large Language Models Are Effective Human Annotation Assistants, But Not Good Independent Annotators

Abstract
Event annotation is important for identifying market changes, monitoring breaking news, and understanding sociological trends. Although expert annotators set the gold standards, human coding is expensive and inefficient. Unlike information extraction experiments that focus on single contexts, we evaluate a holistic workflow that removes irrelevant documents, merges documents about the same event, and annotates the events. Although llm-based automated annotations are better than traditional tf-idf-based methods or Event Set Curation, they are still not reliable annotators compared to human experts. However, adding llms to assist experts for Event Set Curation can reduce the time and mental effort required for Variable Annotation. When using llms to extract event variables to assist expert annotators, they agree more with the extracted variables than fully automated llms for annotation.
[…]
?reference?
Large Language Models Are Effective Human Annotation Assistants, But Not Good Independent Annotators
Feng Gu     Zongxia Li     Carlos Rafael Colon Benjamin Evans     Ishani Mondal     Jordan Lee Boyd-Graber University of Maryland, College Park {fgu1, zli12321, raffy, benevans, imondal, jbg}@umd.edu
1 Introduction
Quality data is crucial for making informed decisions. Bloomberg [^1] https://www.bloomberg.com/, for example, automatically extracts top-quality data from trusted sources. Companies like Scale AI [^2] https://scale.com/ collect, curate, and annotate data for ai models (OpenAI) and applications (self-driving cars, mapping, AR/VR, robotics). [^3] https://openai.com/
While human annotators excel at reference resolution and aligning extracted information, machine coding can handle larger inputs, enhancing data collection efficiency and scalability Mason et al. (2012); King and Lowe (2003). The Uppsala Conflict Data Project Georeferenced Event Data, one of the earliest event datasets relying on human labour, covers only fewer than 300,000 events in over twenty years Sundberg (2013). In comparison, automated data processes like the gdelt Project collect trillions of datapoints The GDELT Project (2021).
Figure 1: Our workflow for annotating events data begins with preprocessing incoming media news. A Support Vector Machine identifies highly relevant documents for manual review. During Event Set Curation, human annotators create unique event sets. Finally, annotators code the domain-specific variables. We apply lm-based similarity indices and use lm-extracted variables to aid manual processing.
Annotating event datasets presents significant challenges. First, statistical models often encounter challenges with text extraction and formatting when processing large corpora automatically Schrodt and Van Brackle (2013). Issues can arise from unreliable source documents, incorrect translations, conflicting information, and inadequate filtering. Second, empirical evaluations yield inconsistent results with extremely low confidence intervals O’Brien (2013). Third, unlike document-level annotation tasks, event annotation requires expertise in synthesizing information from multiple documents Gao et al. (2024); Li et al. (2024a), which often involves additional steps beyond labelling, extraction, or annotation.
Contemporary event datasets typically combine quantitative methods with manual efforts. Algorithms like unsupervised clustering group documents  Golub (2006); Hakim et al. (2014); Probierz et al. (2022); Bouras and Tsogkas (2013); Li et al. (2025a) before human annotators refine the results and extract information Elliott (2018); Pierre and Jackson (2014).
We present a case study on ai automation for gtd  [^4] https://www.start.umd.edu/data-tools/GTD (Global Terrorism Database), whose researchers maintain a similar open-source terrorism database used by various organizations. By integrating ai into the event annotation pipeline, we demonstrate that state-of-the-art nlp techniques can reduce human effort in traditional workflows. Specifically, we show that large language models (llms) are far from expert quality annotations, but they can serve as effective assistants for humans to extract event variables, thus reducing the variable coding time of experts.
More broadly, we outline the potential of integrating llms into event annotation pipelines. Capturing semantic similarity and coding diverse variables with high expert agreement, lms show strong promise for real-world annotation tasks.
2 Related Work
Single-document Event Extraction: Prior work on information extraction focuses primarily on sentence and document-level tasks like part-of-speech tagging Roth and Zelenko (1998) and named-entity recognition Collobert et al. (2011), creating various benchmarks. Gao et al. (2024) introduces cross-document event extraction, which involves merging duplicates and resolving conflicts. Our task aligns with cross-document event extraction but is more complex: an lm must identify documents about the same event and extract information adhering to a strictly defined schema.
Variable Extraction as a qa Task: Neural network models excel in qa benchmarks Hermann et al. (2015); Li et al. (2025b). Extracting domain-specific variables can be framed as a qa task, leveraging contextual information from coding criteria. However, evaluating qa systems is challenging in contextual or complex output spaces Xu et al. (2023); Bulian et al. (2022), often lacking ecological validity. We use trained rule-based and neural matching methods to achieve a more robust variable equivalence detection.
Text Segmentation: Text segmentation divides large text blocks into smaller, topically coherent sections. Koshorek et al. (2018) create an annotated corpus using Wikipedia pages, while Aumiller et al. (2021) use a bert-based model for legal documents. Neural models have achieved excellent performance Glavaš and Somasundaran (2020). Although algorithms like TextTiling Hearst (1997) and TopicTiling Riedl and Biemann (2012) can handle texts with distinct sections, documents with closely related events like ours pose greater segmentation challenges.
llm-assisted Annotation: Annotating domain-specific data requires expertise. Recent nlp advancements enable llms to annotate data efficiently, achieving high accuracy on document-level datasets Tan et al. (2024). However, Tseng et al. (2024) show lms cannot replace experts in financial data annotation. Similarly, Yadav et al. (2024) finds gpt-4 performs poorly in annotating word semantics. Our work explores llm-assisted annotation in a real-world even annotation workflow.
3 A Case Study of Event Annotation
The existing data workflow at the gtd consists of automated and manual processing (Figure 1). To obtain quality data from news articles that can be later used for analysis, experts 1) compile a list of unique events from a collection of documents and 2) extract synthesized event information.
3.1 Event Set Curation
An event set is a collection of documents about the same event. Unlike document-level annotations, event details are scattered across multiple documents, each offering partial or complementary information. Annotators must prioritize informative documents (e.g., those with high tf-idf scores), systematically review them to identify unique events, and create event sets that meet predefined criteria and sending them for downstream coding. For example, one document describes a gun attack “in San Carlos City on Sunday, where armed men opened fire on a vehicle in Barangay Palampas,” while another specifies that “the shooting occurred in Barangay Palampas, San Carlos City, Negros Occidental, on February 20,” providing precise details about the incident’s date.
To ensure consistency in event sets for downstream comparisons, we evaluate candidate event sets using precision, recall, and $F_{1}$ score.
3.2 Variable Coding
Annotators synthesize information from documents in event set for each event, identifying variables like location and perpetrators [^5] Full list and description in Appendix A.5.. One document describes a violent incident involving “approximately 20 people, some armed with axes, who targeted security guards, workers, and police officers” in “Houston, British Columbia,” while another document states that “far-left anti-pipeline extremists” attacked the “Morice River drill pad site off the Marten Forest Service Road.” Annotators use information from all relevant documents to code variables. These variables include text, enumerated types, and integers, representing diverse data types. We measure coding time with and without llm-extracted variables and calculate the frequency of annotators choosing llm-extracted variables.
4 Efficient Event Curation with embedding and llm
Real-world data like news digest often describe multiple events in the same document, posing challenges for tf-idf, which struggle to capture semantic relationships between texts Bengio et al. (2003); Mikolov et al. (2013). We compare pre-trained transformer embeddings and an llm with tf-idf.
4.1 Methodology
We randomly sample $500$ of $7{,}941$ documents from February 2022, unseen by expert annotators, to evaluate similarity. We compare automatically create event sets against the expert-curated reference set. We describe our algorithms in Appendix A.3 and prompts in A.7.
embedding: We use baai Embeddings Xiao et al. (2023) to calculate pairwise similarity, employing grid search to find the threshold maximizing the $F_{1}$ score. Articles with similarity above this threshold (0.8) are clustered.
llm-cls: We use gpt-4o-mini Achiam et al. (2023) to determine if two documents describe the same event.
llm-cls+seg: To address multiple events, we prompt gpt-4o-mini to segment documents before applying llm-cls for binary prediction (Appendix A.7).
4.2 Results and Discussion
Precision Recall $F_{1}$ Identical Sets
tf-idf 0.19 0.09 0.10 12
embedding 0.89 0.51 0.59 66
lm-cls 0.36 0.35 0.35 105
lm-cls+seg 0.65 0.66 0.63 203
Table 1: In Event Set Curation, embedding has the highest precision of creating event sets at 0.89. lm-cls+seg demonstrates superior recall and a higher overall $F_{1}$ score. Additionally, compared to 371 annotated event sets, lm-cls+seg generates event sets that align most closely with expert annotations, nearly doubling the amount of the second-best method, lm-cls.
embedding captures semantic relationships, significantly improving performance with $89\%$ precision. Unlike tf-idf, which attaches less than $10\%$ of relevant documents and has less than $20\%$ relevance among attached documents, embedding reduces false negatives by $82\%$ and identifies about half of the relevant documents. Overall, embedding has a $0.59$ $F_{1}$ score (Table 1).
llm Segmentation helps. Dividing documents into multiple events improves event set curation, yielding the best overall $F_{1}$ score (Table 1).
5 llms Help Experts Code Variables
Human lm Overlap Average
Manual 271 (160) 233 (227) 220 (145) 236 (169)
Hybrid 260 (185) 175 (202) 173 (132) 197 (169)
Average 265 (174) 203 (216) 197 (140) 217 (170)
Table 2: The average time, in seconds along with standard deviations, taken by human annotators to code variables. Annotators consistently spend less time when they have access to lm-extracted variables. When annotators and lm-cls+seg concur on event sets, they take the least amount of time to annotate these sets.
Although embedding has the highest precision, recall is more important because missing document can significantly bias the downstream coding task. llm-cls+seg not only has the highest recall but also finds many event sets identical to expert annotations ($203$ out of $371$), outperforming llm-cls ($105$), embedding ($66$), and tf-idf ($12$) (Table 1). While llm-cls+seg can generate identical event sets, it may also produce problematic ones, negatively impacting annotation. To assess the effect of incorrect event sets in Variable Coding, we replace some expert-curated events with their nearest neighbours in llm-cls+seg, creating three evaluation settings:
1) Human: events annotated by humans but absent in llm-cls+seg.
2) lm: events in llm-cls+seg but not in human-annotated sets.
3) Overlap: events present in both human-annotated and llm-cls+seg sets.
5.1 Methodology
We divide the event sets into three settings:
Manual: Annotators code the variables without llm assistance, serving as the control group.
Automated: We prompt gpt-4o-mini to extract variables from the documents.
Hybrid: Annotators see the llm-extracted variables and can pre-populate them.
Human annotators may prefer llm-extracted variables because seeing prepopulated llm variables before manually entering them can cause automation bias An et al. (2023). To ensure robust result, we duplicate a subset of events and assign them to three expert teams for calculating inter-human agreement. We compare this inter-human agreement with the agreement between the manual and the automated settings (human–lm in Figure 2) to assess whether llm-extracted variables positively influence variable selection frequency. Each team codes $212$ event sets. For each event, annotators code nine variables.
Traditional metrics can be overly strict Kocmi et al. (2021); Chen et al. (2020), and we use two additional equivalence metrics besides exact match:
1) Normalized Match (nm) checks if two variables are identical after string normalization– removing articles and punctuations.
2) bert, fine-tuned on human labeled and synthetic datasets, measures embedding similarity Bulian et al. (2022).
3) pedants uses optimized learned $F_{1}$ and tf-idf encoding to measure matching between variables Li et al. (2024b).
5.2 Results and Discussion
Variable Frequency(%) Count
Country 92.1 278
Location 40.1 278
Target 64.5 265
Perpetrator 78.5 209
Generic Attack 58.5 241
Generic Weapon 73.7 218
Specific Weapon 61.3 194
Kills 79.4 106
Wounds 67.8 97
Overall 66.0 1886
Table 3: Total number of non-NA variables and how often they are chosen by human annotators in the Optional setting. Annotators select extracted variables 66% of the time on average. Notably, the variable Country has a high selection frequency of over 92%. In contrast, annotators choose Location only about 40% of the time.
llm-extracted variables reduce annotation time. Experts code fastest when their judgment align with llm-cls+seg on event sets (Table 2). Coding event sets from human annotators takes the longest. Despite errors in llm-extracted variables, extracted variables reduce annotation time by $25\%$.
Figure 2: The agreement between the manual and the automated settings is comparable. On average, annotators and lm agree 50% using nm. pedants and bert show higher agreements. The difference between human-human and human-lm agreement is not statistically significant, suggesting that the lm-extracted variables provide approximately human-level utility.
Annotators align with lm-extracted variables. Annotators agree with them over half of the time even without seeing them (Figure 2). Human-lm agreement nearly matches human-human agreement, indicating that llm-extracted variables approximate human-level utility.
Figure 3: Agreement by event set type and setting. Annotators show higher agreement in the hybrid setting, where extracted variables are available. This indicates that these variables help code the events. Furthermore, the extracted variables prove particularly beneficial in lm-generated incident sets, which often contain misinformation.
Annotators use llm-extracted variables in the hybrid setting. Annotators adopt extracted variables $66\%$ of the time on average (Table 3). The most selected variable, Country, is chosen over $90\%$ of of the time, while the least selected, Location, is still useful $40\%$ of the time. The hybrid setting shows higher agreement than the manual setting across all event set types (Figure 3), suggesting llm-extracted help coding.
6 Error Analysis
In Event Set Curation, automatically generated event sets correlate poorly with expert annotations. Variable Coding face issues from interpretable errors and schema errors.
Automatic Event Set Curation: While being significantly better than the current baseline, llm-cls+seg still includes over $30\%$ of false positives and negatives, falling short of expert annotators. Nonetheless, embedding and llm-cls+seg prove useful when document tf-idf scores are identical.
Underspecified Instructions: Unclear guidelines lead to imprecise coding by the llm. Table 3 shows high selection frequency for Country but low for Location. While Country is specific, Location is vague, ranging from broad regions to specific sites. For instance, a variable might refer to the administrative area “Pale, Sagaing” in Myanmar, while annotators prefer the specific “Einmahti village”.
Superfluous Information: Documents often contain multiple events, resulting in conflicting reports and causing llm to extract data for the wrong event or get contradictory variables. For example, reports on a Colombian strike vary, with some stating “at least eight people wounded, three seriously,” and others report “five people injured”.
Interpretive Result: Human judgment influences coding. The Facility/Infrastructure category in Generic Attack refers to non-human entities. If buildings are the primary target, it is coded as Facility/Infrastructure; otherwise, it is Armed Assault. Human annotators may disagree on primary targets.
7 Conclusion
Current event annotation processes depends significantly on human judgment. Recent advancements in nlp offer promising potentials for downstream tasks. In this study, we use an llm to reduce human effort in Event Set Curation and Variable Coding. Our results show that an llm outperforms traditional baselines in Event Set Curation. For Variable Annotation, the existing workflow benefits from reduced annotation time. Although fully relying on SoTA llms is still far from expert quality annotations, using an out-of-the-box llm proves useful in helping experts with diverse variable types in a realistic task.
8 Limitations
nlp techniques have not yet reached human-level accuracy in document classification. In Event Set Curation, the computational cost of pairwise similarity increases quadratically with the linear growth in the number of documents, rendering lm-based methods inefficient for large document sets. Although it is impractical to fully replace tf-idf, embedding finds semantically similar documents when tf-idf are the same. To alleviate computational efforts, instead of computing pairwise embedding similarity for every document, we only compute pairwise embedding similarity for documents above the tf-idf threshold. For Variable Coding, extracting variables requires greater granularity to meet the specific criteria outlined in the coding guidelines. These steps are essential for implementing lm-assisted large-scale event data collection workflows.
Appendix A Appendix
Figure 4: Agreement grouped by variable type. Human annotators agree more with extracted variables with higher degree of specificity. Country has over 90% agreement. Generic attack type and weapon type also high agreement. In comparison, low specificity variables like location demonstrate low agreement with human judgment.
A.1 Automated Data Processing
A retrieval model collects news articles from trusted sources. String filters identify documents containing potential terrorism attack information, focusing on keywords such as assault, hostage, and rebel. Then, an algorithm removes duplicates and irrelevant documents from this pool. A Support Vector Machine (svm) uses tf-idf to flag highly relevant documents for manual review.
A.2 Definition of Terrorism and Inclusion Criteria
The cuss team defines a terrorist attack as the threatened or actual use of illegal force and violence by a non-state actor to attain a political, economic, religious, or social goal through fear, coercion, or intimidation. Specifically, an article must have all of the following attributes to be included in the database: {enumerate*}
The event must be intentional – the result of a conscious calculation on the part of a perpetrator.
The event must involve violence - against either property or people.
The perpetrators of the events must be sub-national actors. State-level terrorism is excluded from the database.
In addition, the article must also met at least two of the following criterion: {enumerate*}
The act must be aimed at attaining a political, economic, religious, or social goal. In terms of economic goals, the exclusive pursuit of profit does not satisfy this criterion. It must involve the pursuit of more profound, systemic economic change.
There must be evidence of an intention to coerce, intimidate, or convey some other message to a larger audience (or audiences) than the immediate victims. It is the act taken as a totality that is considered, irrespective if every individual involved in carrying out the act was aware of this intention. As long as any of the planners or decision-makers behind the attack intended to coerce, intimidate or publicize, the intentionality criterion is met.
The action must be outside the context of legitimate warfare activities. That is, the act must be outside the parameters permitted by international humanitarian law, insofar as it targets non-combatants.
A.3 Algorithm for Finding the Best Embedding Threshold
Algorithm 1 Embedding Algorithm
1:Input: list of articles
2:Output: best precision, recall, $F_{1}$
3:best precision, recall, $F_{1}$ $\leftarrow 0,0,0$
4:for $i=1$ to steps do
5:     threshold $\leftarrow$ min +(max - $+\frac{\text{(max - min)}\cdot i}{\text{steps}}$
6:     for each article1 in all articles do
[…]
8:         for each article2 in all articles do
9:              calculate similarity
[…]
11:                  similars.add(article2)
12:              end if
13:         end for
14:         if $F_{1}(\text{ref},\text{similars})>\text{best $F_{1}$}$ then
15:              update best $F_{1}$
16:         end if
17:     end for
18:end for
19:return best (precision, recall, $F_{1}$)
A.4 Over- and Under-generation of lm-cls
Using an lm to generate event set candidates would almost always create different numbers of event sets compared to tf-idf, making direct comparison diffucult. To address potential over- or under-generation, we formulate this as a linear assignment problem, optimizing the average $F_{1}$ score between the results and the human-coded reference set (Appendix A.4).
The linear assignment problem is an optimization problem. The objective is to assign a gold event set to a generated event set in such a way that the overall cost is minimized. $\mathcal{C}_{i,j}$ represents the cost of matching event set $i$ in the gold set with event set $j$ in the prediction set. Formally, the optimal assignment has cost
$\sum_{i}\sum_{j}\mathcal{C}_{i,j},$
where
$\mathcal{C}_{i,j}=-F_{1}{(i,j)}$
A.5 Variable Coding Schema
Here, we describe the variables in the annotation process, specified by cuss.
  1. 1.
    Country: the country in which the terrorism event occurred.
  2. 2.
    Location: the most specific location (e.g., village name) in which the terrorism event occurred.
  3. 3.
    Target: the targeted group of the terrorism event.
  4. 4.
    Perpetrator: the group carrying out the terrorism event.
  5. 5.
    Generic Attack Type: One or more of Facility/Infrastructure Attack, Armed Assault, Assassination, Bombing/Explosion, Hostage Taking (Kidnapping), and NA.
  6. 6.
    Generic Weapon: One or more of Explosives, Firearms, Incendiary, Sabotage Equipment, Melee, Vehicle, and NA,
  7. 7.
    Specific Weapon: A detailed description of Generic Weapon.
  8. 8.
    Kills: Number of people killed during the terrorism event.
  9. 9.
    Wounds: Number of people injured during the terrorism event.
A.6 Human-lm Agreement by Variable Type
Figure 4 shows how often human annotators agree with lm-extracted variables using equivalence metrics. Annotators show 0.89 agreement with Country, a variable with high degree of specificity. In contrast, annotators agree with Location infrequently, suggesting less utility of variables with lower degree of specificity.
A.7 gpt-4o-mini Prompts
The specific model we used is gpt-4o-mini-2024-07-18 with zero temperature. Output might change for future models.   Prompt for document classification:
"Determine whether the following articles describe the same incident:
{article 1}
{article 2}"
  Prompt for Document segmentation:
"The following document describes zero or more terrorism incidents. Segment the document based on incidents mentioned and return an array.
{article}"